### v1

# iMAD: Intelligent Multi-Agent Debate for Efficient and Accurate LLM Inference

Abstract
Large Language Model (LLM) agent systems have advanced rapidly, driven by their strong generalization in zero-shot settings. To further enhance reasoning and accuracy on complex tasks, Multi-Agent Debate (MAD) has emerged as a promising framework that engages multiple LLM agents in structured debates to encourage diverse reasoning. However, triggering MAD for every query is inefficient, as it incurs substantial computational (token) cost and may even degrade accuracy by overturning correct single-agent answers. To address these limitations, we propose intelligent Multi-Agent Debate (iMAD), a token-efficient framework that selectively triggers MAD only when it is likely to be beneficial (i.e., correcting an initially wrong answer). To achieve this goal, iMAD learns generalizable model behaviors to make accurate debate decisions. Specifically, iMAD first prompts a single agent to produce a structured self-critique response, from which we extract 41 interpretable linguistic and semantic features capturing hesitation cues. Then, iMAD uses a lightweight debate-decision classifier, trained using our proposed FocusCal loss, to determine whether to trigger MAD, enabling robust debate decisions without test dataset-specific tuning. Through extensive experiments using six (visual) question answering datasets against five competitive baselines, we have shown that iMAD significantly reduces token usage (by up to 92%) while also improving final answer accuracy (by up to 13.5%).
Figure 1: Overall workflow of iMAD.
1 Introduction
With the rapid progress in Large Language Models (LLMs), agent systems have shown impressive zero-shot reasoning capabilities across tasks such as (visual) question answering, problem solving, or code generation. This ability in zero-shot settings, without access to evaluation data, makes LLM agents appealing for real-world applications by enabling fast and scalable deployment across diverse domains (wan-etal-2023-better; verbalizedconfidencescoresllms). These agent systems typically rely on a single LLM agent to generate step-by-step reasoning, using methods like Chain-of-Thought (CoT) (wei2022chain) or Self-Consistency (wang2022self). However, these approaches often suffer from limited diversity and may overlook alternative reasoning paths as they rely on a single agent’s perspective.
To address this limitation, recent studies have explored multi-agent systems that enable collaboration or debate among multiple agents to enhance reasoning and decision-making (liang-etal-2024-encouraging; chen2023reconcile). Among these approaches, Multi-Agent Debate (MAD), inspired by the role of structured discourse in human cognition, has emerged as a particularly promising framework (wu2023autogenenablingnextgenllm). In MAD, multiple agents independently reason over a query and critique each other’s output through structured interactions, stimulating adversarial dialogue and iterative refinement (liang-etal-2024-encouraging; tillmann2025literaturereviewmultiagentdebate). Such interactions encourage diverse reasoning paths and perspective shifts, enabling agents to recover from faulty initial answers and thus often outperform single-agent systems (liang-etal-2024-encouraging; hsu2025groupthinkmultipleconcurrent).
Despite this benefit, MAD frameworks face two critical limitations that hinder their practical deployment. First, MAD incurs significantly higher computational costs due to iterative LLM queries–measured in total token usage, which includes both input tokens (prompts to each agent) and output tokens (generated responses). Most MAD systems consume three to five times more tokens than single-agent baselines (liu2024groupdebateenhancingefficiencymultiagent), making them costly to scale (see Insight 3). Second, perhaps more counterintuitively, MAD does not consistently improve response quality. Both prior work (zhang2025stopovervaluingmultiagentdebate) and our own empirical analysis show that in many cases, the single-agent output is already correct, making it redundant to trigger MAD. In other cases, the error in the single-agent response cannot be corrected by MAD, or even worse, triggering MAD may override a correct answer with an incorrect one, resulting in degradation in accuracy (see Insight 3). These observations suggest that while MAD can be beneficial, applying it to every query may be inefficient. Not only does it incur substantial computational costs, but it can also degrade accuracy. Therefore, we need an intelligent and principled mechanism that can selectively trigger MAD only when it is likely to be beneficial. This raises a key question: When should a debate be performed to preserve the benefits of MAD while avoiding unnecessary token costs and potential accuracy degradation?
A straightforward approach is to use the confidence score (typically computed as the average log-probability of the output tokens) to estimate single-agent answer correctness. Leveraging this idea, a recent selective MAD framework, called DOWN, attempts to trigger MAD when the confidence score falls below a threshold (eo2025debatenecessaryadaptivemultiagent). However, it requires a threshold tuned on a subset of the evaluation data, which violates the core zero-shot setting assumption. Moreover, even with a fixed threshold, our empirical analysis reveals that confidence scores alone are not reliable indicators of whether MAD is necessary (see Insight 3). We observe that confidence scores could be high even for incorrect answers, revealing the model’s overconfidence. Also, confidence scores are often misaligned with reasoning uncertainty: responses that contain hesitation cues (such as hedging, contradictions, or shallow reasoning) may still receive a high score. This misalignment causes two undesirable situations: skipping the necessary MAD or triggering it unnecessarily.
While existing methods like repeated sampling (i.e., querying the same agent multiple times) or invoking more LLM agents can estimate whether MAD is necessary, they incur high token costs and thus are not scalable (wang2022self). This highlights the need for token-efficient mechanisms that can make informed decisions about when to trigger MAD based on the initial single-agent output. There are two key challenges to achieving this goal:
(C1) How to design an effective prompt that guides the single-agent response to expose richer features for making a debate decision? This decision must rely on features embedded in the initial single-agent output. These features need to be informative and can be efficiently extracted (i.e., without requiring repeated sampling or additional LLM queries).
(C2) How to design a mechanism that intelligently decides when to trigger MAD for efficient and accurate inference in the zero-shot setting? We identify when MAD is more likely to recover incorrect answers without relying on access to the evaluation dataset. Achieving this requires learning generalizable model behavior, which involves addressing the aforementioned issues of LLM agents: overconfidence in incorrect answers and misalignment between confidence scores and semantic cues of hesitation in the response.
To address these challenges, we propose intelligent Multi-Agent Debate (iMAD), a lightweight debate-triggering framework for token-efficient MAD in the zero-shot setting that selectively triggers MAD only when it is more likely to improve the final answer. The overall workflow of iMAD is illustrated in Fig. 1. We summarize our main contributions along with the key components of iMAD as follows:
[…]
  • We evaluate iMAD on three question answering (QA) and three visual question answering (VQA) datasets against five competitive baselines (including two single-agent and three multi-agent frameworks). iMAD reduces token usage by up to 92% while improving accuracy by up to 13.5% through selectively skipping debates that are unnecessary or detrimental. Notably, we train the classifier solely on two representative datasets selected to capture diverse model reasoning behaviors, allowing the classifier to learn generalizable model behavior and perform effectively across six held-out datasets.
2 Related Work
We categorize highly related works into three groups.
Single-Agent and Multi-Agent LLMs.
LLMs have demonstrated strong reasoning capabilities in single-agent settings, where a single LLM agent performs all reasoning steps independently. Foundational methods like CoT prompt models to generate intermediate reasoning steps, enabling better handling of complex tasks (wei2022chain). Recently, Self-Consistency further improves accuracy over CoT by sampling multiple outputs and selecting the most frequent answer (wang2022self; wang2024soft; li2024escape). However, single-agent reasoning relies solely on internal sampling, lacking perspective diversity and explicit self-correction. To address this, multi-agent LLM frameworks leverage multiple agents to reason independently or coordinate through structured interaction, thus improving accuracy over single-agent approaches. Some methods generate multiple outputs for joint evaluation (e.g., CoMM (Chen2024CoMMCM)), while others assign hierarchical agent roles (e.g., Mixture-of-Agents (MoA) to progressively refine reasoning (wang2024moa; chen2024moa; li2025rethinking)). While these methods stimulate collaborative reasoning among agents, they often incur high computational costs and deliver inconsistent improvements over single-agent baselines (zhang2025stopovervaluingmultiagentdebate; cemri2025multi).
MAD Frameworks.
MAD frameworks represent a structured subclass of multi-agent LLM where agents engage in explicit argumentative exchanges (e.g., critiques, rebuttals, or deliberation) to refine initial outputs (tillmann2025literaturereviewmultiagentdebate). Existing MAD methods include role-based debates with assigned affirmative, negative, and moderator roles (liang-etal-2024-encouraging; wang2024unleashingemergentcognitivesynergy), implicit debate via input perturbation and aggregation (e.g., Reconcile (chen2023reconcile)), and intra-agent self-refinement (srivastava2025debate; zhang2024can). Building on these MAD designs, GroupDebate extends MAD by coordinating subgroup debate (liu2024groupdebateenhancingefficiencymultiagent). While MAD enhances interpretability and error correction, recent studies show that it can also introduce noise or overturn correct single-agent answers, thus degrading accuracy (zhang2025stopovervaluingmultiagentdebate).
Confidence-based Selective Debate.
In a concurrent work, the proposed DOWN framework aims to reduce the token cost of MAD by using LLM-generated confidence scores to determine when to trigger debate (eo2025debatenecessaryadaptivemultiagent). However, selecting an appropriate confidence threshold requires labeled data from the evaluation dataset, which violates the zero-shot setting assumption normally upheld by current single-agent and MAD baselines and limits applicability in real-world scenarios. Moreover, confidence scores alone are not reliable in determining whether triggering a debate will improve the answer (see Insight 3). Although DOWN can reduce token usage by skipping debate on high-confidence responses, it often fails to identify cases where debate is beneficial and misses opportunities to correct recoverable errors.
3 Key Insights
Dataset
Single-Agent
Multi-Agent
Debate (MAD)
Acc (%) # Token Acc (%) # Token
MEDQA 76.6 653 81.9 4,034
MMLU 86.8 764 89.5 3,348
GSM8K 71.3 618 76.4 3,446
OKVQA 88.3 1,945 89.8 7,803
VQA-v2 77.5 2,245 81.0 8,796
ScienceQA 86.0 1,720 89.4 6,777
Table 1: Comparison of accuracy (Acc) and average token costs (# Token) per question between single-agent CoT and MAD frameworks across QA and VQA datasets.
While MAD shows promise in enhancing LLM reasoning, its deployment remains limited, as modest accuracy gains often come with substantial computational cost. In this section, we first quantify the token overhead of MAD compared to single-agent CoT. Then, we analyze when MAD is beneficial and reveal that the positive impact exists only for a subset of instances. Finally, we examine whether standard uncertainty heuristics (e.g., based on confidence score) can effectively guide debate triggering.
(Insight 1) MAD achieves a higher accuracy at the cost of a substantial token overhead.
We quantify the trade-off between accuracy and token costs by comparing CoT (wei2022chain) and MAD (liang-etal-2024-encouraging) across six QA and VQA datasets in Table 1. Consistent with prior findings (liu2024groupdebateenhancingefficiencymultiagent), we observe that MAD achieves a higher accuracy than CoT, with gains ranging from 1.5% (on OKVQA) to 5.3% (on MEDQA). However, MAD consumes three to five times more tokens than CoT, primarily due to routing the same query to multiple agents, each requiring separate input prompts and generating individual responses (chen2024optima; eo2025debatenecessaryadaptivemultiagent). This cost is more pronounced in VQA tasks, where visual inputs further increase token usage. These observations indicate that the accuracy gains come at a high computational cost, making MAD impractical to deploy at scale.
(Insight 2) The accuracy gains in MAD are primarily driven by a subset of cases.
While a concurrent work has noted a similar observation (zhang2025stopovervaluingmultiagentdebate), we provide a systematic breakdown to quantify the specific sources of MAD’s accuracy gains. Specifically, we categorize each input instance into four cases: (i) incorrect in single-agent but correct in MAD ($\rightarrow$); (ii) correct in single-agent but incorrect in MAD ($\rightarrow$); (iii) correct in both ($\rightarrow$); and (iv) incorrect in both ($\rightarrow$). As shown in Table 2, the ideal scenario where MAD makes corrections ($\rightarrow$) accounts for a small portion (e.g., 4.9% in OKVQA to 19.1% in GSM8K). In contrast, many debates are either redundant (i.e., single-agent answers are already correct: $\rightarrow$), ineffective (i.e., unresolved single-agent errors: $\rightarrow$), or harmful (i.e., flipping correct single-agent answers to incorrect: $\rightarrow$). This shows that while MAD improves the overall accuracy, the benefit is limited to a small portion of cases. Thus, indiscriminately applying MAD to all cases could waste computational resources and even degrade accuracy.
Dataset $\rightarrow$ (%) $\rightarrow$ (%) $\rightarrow$ (%) $\rightarrow$ (%)
MEDQA 11.9 6.6 70.0 11.5
MMLU 6.9 4.2 82.6 6.3
GSM8K 19.1 14.0 57.3 9.6
OKVQA 4.9 3.4 84.9 6.8
VQA-v2 9.2 5.7 71.8 13.3
ScienceQA 7.9 4.5 81.5 6.1
Table 2: Breakdown of MAD outcomes across datasets: percentage of cases where the MAD flipped the answer correctly ($\rightarrow$) or wrong ($\rightarrow$) from single-agent CoT answer, and percentage of cases where both MAD and single-agent CoT are correct ($\rightarrow$) or wrong ($\rightarrow$).
(Insight 3) Confidence scores are not reliable indicators of whether debate is beneficial or not.
We investigate whether LLM-generated confidence scores can effectively indicate when MAD is likely to improve answers (confidence; eo2025debatenecessaryadaptivemultiagent). A natural hypothesis is that answers with a high confidence score may not need MAD; only answers with a low confidence score could benefit from MAD. To test this, we consider confidence scores from various prompting strategies (confidence; eo2025debatenecessaryadaptivemultiagent) and analyze their alignment with cases where MAD corrects initially incorrect single-agent answers. The Cumulative Density Function (CDF) of confidence scores is presented in Fig. 2. We observe that the CDF is highly right-skewed and poorly aligned with answer correctness or debate effectiveness. Notably, incorrect answers often receive high confidence scores, sometimes even exceeding correct answers, indicating a strong bias of overconfidence. Moreover, even many hesitant or shallow responses still receive inflated confidence scores. This misalignment between the confidence score and the uncertainty expressed in reasoning undermines the effectiveness of heuristics that make debate decisions based on confidence scores only.
Figure 2: Cumulative Density Function (CDF) of confidence scores for correct and incorrect single-agent answers.
4 Our Design: iMAD
In this section, we present iMAD, a token-efficient framework that selectively triggers MAD only when it is likely to correct an initially incorrect single-agent answer. Leveraging the aforementioned insights, iMAD aims to substantially reduce the token overhead of MAD while retaining or even improving the accuracy (Insight 3). We begin with an overview of the iMAD framework (Section 4.1), which integrates structured self-critique prompting with a lightweight classifier to decide whether debate should be triggered (Insight 3). The classifier leverages interpretable features extracted from the single-agent response to assess the need for debate, enabling robust debate decisions in zero-shot settings without dataset-specific tuning. To train this classifier effectively, we propose FocusCal loss ($L_{\text{FC}}$) (Section 4.2) to address the aforementioned issues of LLM: overconfidence and misalignment between confidence scores and semantic uncertainty (Insight 3). This design enables iMAD to make token-efficient and accurate debate decisions based on single-agent responses in zero-shot settings.
4.1 Framework Overview
[…]
Given an input question, the system first prompts the LLM to generate a structured response with three key components: (i) an initial CoT justification supporting the original answer; (ii) a required self-critique presenting a counterargument; and (iii) a final reflection including the chosen answer and explicit confidence scores for both initial reasoning and the self-critique perspectives. This structure stimulates a mini-debate: if both perspectives provide similarly strong or weak reasoning with comparable confidence scores, the model is likely to show internal hesitation, suggesting that MAD could be beneficial. Conversely, if one side presents a clear and well-supported argument while the other is weak, the answer is likely already determined, either confidently correct or confidently incorrect. In the latter case, the internally coherent but flawed reasoning makes it difficult to correct through debate.
[…]
We extract 41 interpretable linguistic and semantic features from the structured single-agent output, drawing from the question, initial reasoning, and self-critique. These interpretable features capture human-understandable cues to reasoning quality and internal hesitation, including surface-level statistics, readability scores, part-of-speech counts (e.g., nouns, verbs, and adjectives), question-type indicators, and lexical cues of uncertainty, such as hedging and contrast. The feature set enables fine-grained detection of subtle uncertainty cues that are often not reflected in the model’s raw confidence scores. These interpretable features help the classifier identify when MAD is likely to be beneficial by capturing uncertainty cues that are not well aligned with confidence scores. Rather than relying on a subset of features, we combine complementary semantic, syntactic, and pragmatic signals to form a holistic view of model behavior. This is crucial for generalization in zero-shot settings without access to an evaluation dataset. A detailed list of all 41 features is provided in the extended version (see Appendix A.2).
[…]
We develop a lightweight MLP-based classifier to decide whether to trigger MAD based on features extracted from a structured single-agent response. The input feature vector $\mathbf{z}\in\mathbb{R}^{d}$, a $d$-dimensional real-valued vector space, where $d$ is the total number of extracted features. The vector $\mathbf{z}$ includes LLM-generated confidence score $p_{\text{LLM}}$, along with semantic and linguistic features from the question, answer, and self-critique. The classifier outputs a scalar $p\in(0,1)$, indicating the likelihood that the single-agent answer is either correct or unrecoverably incorrect (where a debate is considered unnecessary).
During training, we run the single-agent pipeline on each example to generate an answer and self-critique, and then assign a binary label $y$ where $y=1$ if the answer matches the ground truth and $y=0$ otherwise. The classifier is trained with these binary correctness labels, but its goal is not to replicate these labels. Instead, it aims to identify debatable cases: answers likely to be incorrect but potentially correctable through MAD. This distinction is crucial because confidently incorrect answers, often resulting from coherent but flawed reasoning, are unlikely recoverable through MAD.
During inference, we apply a decision threshold $\tau\in(0,1)$, tuned on a validation set, to the predicted score $p$: if $p<\tau$, MAD is triggered; otherwise, the original single-agent answer is retained. A high $p$ indicates triggering MAD is unnecessary, while a low $p$ flags uncertain and potentially recoverable errors that warrant MAD. This selective mechanism enables iMAD to devote computational resources to the most uncertain and error-prone cases, reducing token costs while improving accuracy. The classifier is trained using our proposed FocusCal loss, detail in Section 4.2.
4.2 Debate Decision Classifier with FocusCal Loss
To train the classifier, we propose FocusCal loss ($L_{\text{FC}}$), a composite objective addressing overconfidence and uncertainty in single-agent confidence scores (Insight 3). It combines three components: (i) Asymmetric Focal Loss ($L_{\text{AF}}$) that targets the overconfidence issue by penalizing confidently incorrect predictions more than correct ones, encouraging the model to remain cautious on borderline cases; (ii) Confidence Penalty ($L_{\text{CP}}$) that aligns the predicted score $p$ with an auxiliary uncertainty score $u\in(0,1)$, derived from semantic hesitation features via an MLP, penalizing overconfident predictions with uncertain reasoning; (iii) Expected Calibration Error (ECE) that regularizes predicted scores for empirical calibration (Nixon_2019_CVPR_Workshops). Together, these components help the classifier detect recoverable errors that merit MAD and avoid unnecessary debate.
To realize this design, the classifier passes the input feature vector $\mathbf{z}$ through a shared MLP feature encoder $f_{\text{e}}(\cdot)$ to produce a high-level representation, which is then fed into two separate output heads: a correctness head $f_{p}(\cdot)$ and a hesitation head $f_{u}(\cdot)$, producing two scalar logits $\ell_{\text{p}}$ and $\ell_{\text{u}}$:
$\ell_{\text{p}}:=f_{p}(f_{\text{e}}(\mathbf{z}))~\text{and}~\ell_{\text{u}}:=f_{u}(f_{\text{e}}(\mathbf{z})).$ (1)
To integrate the scalar LLM confidence score $p_{\text{LLM}}$ with the MLP-produced logit $\ell_{\text{p}}$ and keep them mathematically consistent, we convert $p_{\text{LLM}}$ into $\ell_{\text{LLM}}$ in the logit space:
$\ell_{\text{LLM}}:=\log\left(\frac{p_{\text{LLM}}}{1-p_{\text{LLM}}}\right).$ (2)
The predicted score $p$ and the uncertainty score $u$ are then calculated as:
$p$ $:=\sigma(w_{1}\cdot\ell_{\text{LLM}}+w_{2}\cdot\ell_{\text{p}}+\epsilon),$ (3)
$u$ $:=\sigma(\ell_{\text{u}}),$ (4)
where $w_{1},w_{2},\epsilon\in\mathbb{R}$ are learnable parameters and $\sigma(\cdot)$ is the sigmoid function. This two-headed MLP separates the predicted score $p$, used for debate decisions, from the auxiliary uncertainty score $u$, which captures hesitation cues in the reasoning path and is supervised via $L_{\text{CP}}$. This design enables the classifier to better identify hesitant and potentially recoverable errors that warrant debate, while skipping MAD for confidently correct answers or unrecoverable errors.
To optimize the classifier for accurate and calibrated debate triggering, we train it using the FocusCal loss:
$L_{\text{FC}}(y,p,u):=L_{\text{AF}}(y,p)+\lambda\cdot L_{\text{CP}}(y,p,u)+\mu\cdot\text{ECE}(y,p).$ (5)
The weights $\lambda$ and $\mu$ are non-negative coefficients tuned via grid search on a held-out validation set to balance the contributions of uncertainty alignment and calibration. We discuss the details of each term in $L_{\text{FC}}$ below.
Asymmetric Focal Loss ($L_{\text{AF}}$).
This component places the strongest penalty on the cases where the classifier assigns a high predicted score $p$ to an answer that is factually incorrect (i.e., labeled $y=0$). These are exactly the instances where MAD should be triggered but was mistakenly skipped. To address this, we adopt an asymmetric focal loss defined as
$L_{\text{AF}}(y,p):=\begin{cases}-\alpha_{1}(1-p)^{\gamma}\log(p),&\text{if }y=1;\\ -\alpha_{0}~p^{\gamma}\log(1-p),&\text{if }y=0,\end{cases}$ (6)
where $\gamma>0$ is a focusing parameter that down-weighs well-classified examples (i.e., when the predicted score $p$ is close to the ground-truth label), and emphasizes incorrect or harder cases. The class-specific weights $\alpha_{1},\alpha_{0}>0$ control the relative emphasis on each class. We typically set $\alpha_{0}>\alpha_{1}$ to attribute a large penalty to cases where the model assigns a high $p$ to incorrect cases ($y=0$), reflecting overconfident decisions that skip needed debate. This loss formulation directly addresses overconfidence by encouraging the classifier to assign a low predicted score $p$ to incorrect single-agent answers. The asymmetric focal loss $L_{\text{AF}}$ is thus designed to emphasize wrong predictions, many of which can be recovered through MAD, thus helping the classifier better identify cases that warrant debate.
Confidence Penalty ($L_{\text{CP}}$).
To further align the model’s predicted score $p$ with its uncertainty score $u$, we introduce a regularization loss term that penalizes misalignment between them. This uncertainty score $u$ reflects the model’s internal hesitation: a high value of $u$ indicates distributed uncertainty across semantic interpretations, while a low value of $u$ indicates peaked, confident predictions. Although $p$ expresses the overall correctness belief, $u$ offers a complementary perspective by capturing the uncertainty or hesitation embedded in the response features. To reconcile these two signals, we define Confidence Penalty as
$L_{\text{CP}}(y,p,u):=\begin{cases}u^{2},&\text{if }y=0\text{ and }p>\tau,\\ (1-u)^{2},&\text{if }y=1\text{ and }p<\tau,\\ 0,&\text{otherwise}.\end{cases}$ (7)
This confidence penalty term penalizes underconfidence for correct answers (when high uncertainty occurs with a low $p$) and overconfidence for incorrect ones (when low uncertainty occurs with a high $p$). By enforcing the consistency between the predicted score $p$ and the uncertainty score $u$, this loss mitigates the misalignment between the predicted score $p$ and the semantic hesitation signals, leading to more reliable debate triggering decisions.
Single-Agent Full-Debate MAD Selective MAD
Dataset CoT SC MAD GD DOWN iMAD
Acc (%) # Token Acc (%) # Token Acc (%) # Token Acc (%) # Token Acc (%) # Token Acc (%) # Token
MEDQA 76.6 653 77.3 3,482 81.9 4,034 80.2 16,832 79.2 1,161 82.0 1,300
MMLU 86.8 764 88.2 3,772 89.5 3,348 82.6 13,216 88.3 901 89.2 1,010
GSM8K 71.3 618 74.5 3,622 76.4 3,446 73.4 15,321 72.6 812 84.8 1,025
OKVQA 88.3 1,945 89.2 11,031 89.8 7,803 87.3 33,932 88.1 2,344 90.3 2,601
VQA-v2 77.5 2,245 77.6 14,013 81.0 8,796 81.3 36,091 78.6 3,262 81.3 3,489
ScienceQA 86.0 1,720 86.2 9,833 89.4 6,777 87.4 26,924 87.0 2,519 90.8 2,893
Table 3: Accuracy (Acc) and average token cost (# Token) per question comparison of iMAD and baselines across datasets.
Expected Calibration Error (ECE).
To improve the reliability of $p$ as a debate-triggering score, we incorporate ECE to encourage its alignment with the empirical correctness. Suppose the dataset contains $N$ instances indexed by $i\in\{1,\dots,N\}$, where each instance $i$ has a predicted score $p^{(i)}\in[0,1]$ and a binary ground-truth label $y^{(i)}\in\{0,1\}$. Based on $p^{(i)}$, We divide the interval $[0,1]$ into $B$ equal-width bins and assign each instance $i$ to a bin $b\in\{1,\dots,B\}$. Let $\mathcal{I}_{b}$ denote the set of indices $i$ whose predicted scores falls into bin $b$. We then compute ECE as:
$\text{ECE}(y,p):=\sum_{b=1}^{B}\frac{1}{N}\left|\sum_{i\in\mathcal{I}_{b}}p^{(i)}-\sum_{i\in\mathcal{I}_{b}}y^{(i)}\right|.$ (8)
For each bin, we measure the average absolute difference between predicted scores and ground-truth labels. Minimizing ECE aligns $p$ with empirical correctness, leading to more reliable decision boundaries for triggering MAD.
5 Evaluation
In this section, we evaluate iMAD across six datasets, measuring both token efficiency (i.e., token usage per question) and accuracy (i.e., final answer correctness). We further analyze the quality of iMAD’s debate decisions on whether to trigger or skip MAD, and how often these decisions lead to beneficial outcomes. Finally, we present ablation studies on the impact of structured self-critique prompting and FocusCal loss in the extended version (see Appendices C.1 and  C.2, respectively).
5.1 Experimental Setup
Datasets. We conduct experiments on six diverse datasets spanning both textual QA and image-text VQA domains. The QA benchmarks include (i) MedQA: USMLE-style medical QA (jin2020disease); (ii) MMLU: professional exam questions (hendrycks2020measuring); and (iii) GSM8K: grade-school math word problems (cobbe2021gsm8k). For VQA, we use (iv) OKVQA: emphasizes visual questions that require external knowledge (okvqa); (v) VQA-v2: natural image QA with reduced bias (balanced_vqa_v2); and (vi) ScienceQA: multimodal science questions from school curricula (lu2022learn).
Baselines.
We compare iMAD with five strong baselines across single-agent, full-debate MAD, and selective MAD approaches. For single-agent methods, we include: (i) CoT (wei2022chain); and (ii) SC (wang2022self), which runs CoT five times and selects the answer via majority voting. For full-debate MAD approaches that trigger debate for all instances, we consider: (iii) MAD (liang-etal-2024-encouraging), using agents with distinct personas in a three-agent setup; and (iv) GD (liu2024groupdebateenhancingefficiencymultiagent), which clusters 5 agents into subgroups for parallel discussion, followed by 3 rounds of inter-group consensus voting. For selective MAD triggering, we evaluate: (v) DOWN (eo2025debatenecessaryadaptivemultiagent). Unlike other baselines, DOWN requires labeled evaluation data to tune its threshold. For fair comparison, we use a threshold of 0.8 as reported most effective in the original paper.
Metrics.
We evaluate each method using two key metrics: (i) total (input + output) token usage per data instance; (ii) answer accuracy, measuring final answer correctness; (iii) accuracy per 100k tokens (ApT), measuring token efficiency (see Appendix B.1); and (iv) per-question inference time, measuring computational efficiency (see Appendix B.2).
Classifier Configurations.
We use a lightweight MLP with six fully connected layers of 200 hidden units each, the batch normalization, ReLU activations, and a dropout rate of 0.2. We train the model for 50 epochs using the Adam optimizer (adam) with learning rate of 0.001 on standardized features via StandardScaler (de2023choice). To support task-agnostic generalization, we train the classifier on PubMedQA (pubmedqa) and GQA (GQA), which are not included in the evaluation. We set the FocusCal loss hyperparameters as $\alpha_{\text{0}}=2.0$, $\alpha_{\text{1}}=1.0$, $\gamma=2$, $\lambda=6$, $\mu=5$, and $B=15$ bins for ECE. The debate threshold is set to $\tau=0.7$. We chose these hyperparameters via grid search on a held-out validation set to ensure stable training and strong performance across all datasets.
Implementation Details.
For LLM-based prompting, we use Gemini 2.0 Flash as the LLM agents for the primary results. We also evaluate GPT-5 nano and Qwen 3.0 (see Appendix C.3 for details and results). By default, we use a temperature of 0.0 to ensure deterministic outputs and a maximum sequence length of 512 tokens. We present the prompt templates details in the extended version (see Appendix A.1). We conducted all MLP training and inference using a single NVIDIA RTX 4090 GPU.
Dataset Skipped (%) Triggered (%)
Good Bad Bad Good
$\!\rightarrow$ $\!\rightarrow$ $\!\rightarrow$ $\!\rightarrow$ $\rightarrow \!\rightarrow \!\rightarrow \!\rightarrow$
MEDQA 10.5 66.3 4.9 4.8 1.0 3.7 1.7 7.1
MMLU 5.8 80.0 3.2 3.5 0.5 2.5 1.1 3.4
GSM8K 6.8 57.1 11.3 2.9 2.7 0.3 2.7 16.2
OKVQA 6.8 83.7 2.8 2.3 0.0 1.2 0.6 2.6
VQA-v2 12.9 69.7 5.3 4.9 0.4 2.1 0.4 4.3
ScienceQA 6.1 78.0 3.6 2.2 0.0 3.5 0.9 5.7
Table 4: Breakdown of iMAD debate decisions by whether MAD was skipped or triggered, and whether the decision was beneficial (Good) or harmful (Bad) based on changes in accuracy or token cost across datasets.
5.2 Main Results
We evaluate the token efficiency and accuracy of iMAD across six diverse datasets, comparing it against three categories of baselines: single-agent methods (CoT and SC), full-debate MAD frameworks (MAD and GD), and a confidence-based selective MAD approach (DOWN).
Advantage over Single-Agent and Full-Debate MAD Baselines.
As shown in Table 3, iMAD achieves superior token efficiency while maintaining or improving accuracy over all single-agent and full-debate MAD baselines. While CoT uses the fewest token, iMAD achieves up to 13.5% higher accuracy. Compared to SC, iMAD drastically reduces token usage with consistently higher accuracy. For example, on MEDQA, iMAD reduces token cost by 62.7% while improving accuracy by 4.7%. Against full-debate MAD baselines, iMAD reduces token usage significantly while achieving comparable or higher accuracy. On MEDQA, iMAD uses 68% fewer tokens than MAD and 92% fewer than GD, while achieving the highest accuracy among the three. A key to iMAD’s zero-shot efficiency is the classifier’s strong generalization: we trained it on two representative datasets selected to capture diverse model reasoning behaviors, enabling it to learn generalizable model behavior. The inefficiency of the full-debate strategy is especially evident on GSM8K, where iMAD outperforms MAD by 8.4% in accuracy. The only exception is MMLU, where iMAD slightly underperforms MAD. As shown in Table 4, we skip 3.5% of necessary debate, slightly higher than the 3.4% where a triggering debate corrects the error. Many questions in MMLU are short, factual, and span diverse domains, so incorrect single-agent answers often sound fluent and highly confident, with few hesitation cues. As a result, the debate classifier is more likely to skip debates that are actually needed.
Advantage over Confidence-Based Selective MAD Triggering.
As shown in Table 3, DOWN and iMAD incur comparable token costs, with DOWN using slightly fewer tokens by omitting self-critique and skipping some needed debates. However, it consistently underperforms both iMAD and full-debate MAD in accuracy as it cannot tune thresholds in the zero-shot setting. In contrast, iMAD incurs a slight token overhead due to self-critique prompting and more necessary debates, but this cost is justified by improved identification of recoverable incorrect answers. For example, on OKVQA, DOWN’s accuracy remains near the single-agent baseline, revealing its inability to identify when debate is truly needed. This limitation arises because DOWN learns data-specific model behavior rather than generalizable behavior. In contrast, iMAD captures hesitation cues through a classifier trained on diverse features, enabling more accurate debate decisions and achieving higher accuracy with fewer tokens in zero-shot settings.
5.3 Breakdown of iMAD Debate Decisions
To evaluate the effectiveness of iMAD’s selective debate triggering, we analyze its decision outcomes across all datasets in Table 4. For each instance, we precompute both the single-agent and MAD outputs to establish whether MAD improves the answer, serving as the ground truth for evaluating decisions. We then measure how often iMAD is aligned with these beneficial outcomes. Overall, up to 95.9% of iMAD’s decisions are beneficial. When skipping debate, iMAD preserves correct answers ($\rightarrow$✓) in 65-80% of cases and avoids wasted computation on unrecoverable errors ($\rightarrow$✗) up to 13%. When triggering debate, iMAD often recover incorrect answers ($\rightarrow$✓). For example, iMAD successfully flips 16.2% of cases on GSM8K and 7.1% on MEDQA, approaching their respective upper bounds of 19.1% and 11.9% (see Table 2). Crucially, harmful decisions, such as overturning correct answers ($\rightarrow$✗) or incurring unnecessary debate overhead ($\rightarrow$✗ and $\rightarrow$✓), remain consistently low (typically around 5-10%). These results highlight iMAD’s ability to selectively trigger MAD only when it is likely to improve accuracy, while avoiding unnecessary token costs.
6 Conclusion
We presented iMAD, a token-efficient MAD framework that selectively triggers debates only when they are likely to improve outcomes. It combines structured self-critique prompting with a lightweight debate-decision classifier trained using our proposed FocusCal loss, enabling effective debate decisions based on single-agent responses without using evaluation data. Compared with five diverse baselines, iMAD reduces token usage by up to 92% while improving accuracy by up to 13.5%. These results show that iMAD is a practical and scalable solution for collaborative reasoning in agentic LLM systems. Future work includes exploring adaptive or online learning approaches to reduce labeling costs during classifier training and further improve generalization, as discussed in the extended version (see Appendix D).
Acknowledgments
This research was supported in part by NSF grant CNS-2315851, the Commonwealth Cyber Initiative (CCI), and a Virginia Tech Presidential Postdoctoral Fellowship.
Appendix
We further include implementation details, additional analysis, additional results, and discussion in the appendix. First, we provide the exact prompt templates for all agents in QA and VQA tasks, including shared context, memory synchronization, role definitions and outputs, structured JSON formats, and placeholder fields in Appendix A. We also summarize all 41 interpretable features used in our debate decision classifier. In Appendix B, we present additional analysis on token efficiency and inference time, and we report ablation studies on the structured self-critique prompt and FocusCal loss, as well as cross-LLM results using Qwen 3.0 and GPT-5 nano in Appendix C. Lastly, we discuss our assumptions, limitations, and future work in Appendix D.
Appendix A Implementation Details
A.1 Prompt Templates for VQA and QA Tasks
This section specifies the full prompt protocol used in our framework for VQA and QA tasks. We first describe how MAD maintains shared context and memory synchronization across agents, the roles and outputs of each agent, the required structured JSON format used by the Judge, and the placeholders that instantiate prompts at runtime. We then list the exact task-specific prompt templates for VQA and QA, specifying for each agent what it receives (inputs), the instruction it follows, and the required output.
Shared Context and Memory Synchronization.
When MAD is triggered, all agents read the same dialogue history. At the beginning of each round, we share the running transcript with every agent, which includes the question, the initial reasoning or self critique, and all prior messages from all agents, so that decisions are made with a shared context. This synchronization keeps the discussion coherent.
Roles and Outputs.
We use three roles in MAD: Debater (Affirmative), Debater (Negative), and Judge, plus a single agent baseline and a self critique stage in iMAD. Debaters must select exactly one option and provide a concise justification grounded in the available evidence (image for VQA, text for QA). The Judge returns structured JSON so that its decision can be parsed automatically, while the self-critique single agent produce free form text that is later featurized for the debate decision classifier.
Structured JSON Output.
At the end of each round, the Judge returns a single JSON object with the fields {"Preference", "Supported Side", "Reason", "debate_answer"}. The system uses this object to control the debate: if "Preference" is "No", we start the next round; if it is "Yes", we stop the debate and record "debate_answer" when provided. For finalization, the Judge returns a JSON output with {"Reason", "debate_answer"}. We reject malformed JSON outputs such as extra text, missing keys, or invalid values and then reissue the prompt.
Placeholders.
We fill each template at runtime with concrete values drawn from the example and the current debate state. Example fields: {QUESTION}, {OPTIONS} (current option set), {OPTIONS2} (optional reduced set for the final decision), and {IMAGE} for VQA. Debate state fields: {ROUND} (current round index), {TRANSCRIPT} (the synchronized dialogue history), {AFF_ANS} and {NEG_ANS} (the latest messages from each side), and {OPP_ANS} (the opponent’s most recent message). Before each agent is prompted, we compute these values from the input example and the live transcript and then substitute them into the template.
Prompt Template for QA Tasks.
We present the QA prompt templates in execution order. The self critique single agent receives the question and the option set, produces an answer with reasons, and provides a concise self critique. When MAD is triggered, the Debater (Affirmative) selects one option and justifies it, and the Debater (Negative) challenges the affirmative position by arguing for a different option with a brief rationale. Both debaters read the synchronized transcript so decisions are made on a shared context. Finally, the Judge returns structured JSON that either requests another round or finalizes the answer by providing debate_answer and a short reason.
Self-Critique Single-Agent Baseline Receives: {QUESTION}, {OPTIONS}. Instruction: Answer the multiple-choice question. Use a step-by-step chain of thought to justify your initial answer. Then, argue against your own answer with a plausible counter-reasoning — this is not optional. Finally, confirm your final answer and report confidence scores for both the initial belief and the forced disagreement. Output: One option (A/B/C/…); a short justification.
Debater (Affirmative) Receives: {TRANSCRIPT}; {QUESTION}, {OPTIONS}; initial stance {AFF_ANS}; and {OPP_ANS} from round 2+. Instruction: First turn — restate {AFF_ANS} and justify. Later turns — agree or refute {OPP_ANS}; if refuting, choose an alternative strictly from {OPTIONS} and justify. Output: A short argument plus one choice from {OPTIONS}.
Debater (Negative) Receives: {TRANSCRIPT}; {QUESTION}, {OPTIONS}; latest affirmative message {AFF_ANS}; {OPP_ANS}. Instruction: Disagree when warranted. Provide an alternative strictly from {OPTIONS} and justify succinctly. Output: A counter-argument plus one choice from {OPTIONS}.
Judge (per round) Receives: {TRANSCRIPT}; {QUESTION}, {OPTIONS}; {AFF_ANS}, {NEG_ANS}. Instruction: Evaluate accuracy and justification quality; determine preference and tentative answer if any. Output (JSON only): {”Preference”:”Yes—No”,”Supported Side”:”Affirmative—Negative”,”Reason”:”…”,”debate_answer”: one of {OPTIONS} or ””}.
Judge (Finalization) Receives: {TRANSCRIPT}; {QUESTION}, {OPTIONS}; candidate list; {OPTIONS2}. Instruction: (A) List candidates strictly from {OPTIONS}. (B) Provide the final answer with a brief reason. Output (JSON): {”Reason”:”…”,”debate_answer”: one of {OPTIONS2}}.
Prompt Template for VQA Tasks.
We present the VQA prompt templates in execution order. The self critique single agent receives the image, the question, and the option set, produces an answer with reasons grounded in visual evidence, and provides a concise self critique. When MAD is triggered, the Debater (Affirmative) selects one option and justifies it using concrete visual cues such as objects, attributes, or regions in the image, and the Debater (Negative) challenges the affirmative position by arguing for a different option with a brief image-grounded rationale. Similar to the QA task, both debaters read the synchronized transcript so decisions are made on a shared context. The Judge follows the same protocol: it returns structured JSON that either requests another round or finalizes the answer by providing debate_answer and a brief reason, with an explicit check for consistency with the image.
Self-Critique Single-Agent Baseline Receives: {IMAGE}, {QUESTION}, {OPTIONS}. Instruction: From the image, answer the multiple-choice question. Use a step-by-step chain of thought to justify your initial answer. Then, argue against your own answer with a plausible counter-reasoning — this is not optional. Finally, confirm your final answer and report confidence scores for both the initial belief and the forced disagreement. Output: One option (A/B/C/…); a short justification.
Debater (Affirmative) Receives: {TRANSCRIPT}; {IMAGE}, {QUESTION}, {OPTIONS}; initial stance {AFF_ANS}; and {OPP_ANS} from round 2+. Instruction: First turn — restate {AFF_ANS} and justify from the image. Later turns — agree or refute {OPP_ANS}; if refuting, select an alternative strictly from {OPTIONS} and justify briefly. Output: A short argument plus one choice from {OPTIONS}.
Debater (Negative) Receives: {TRANSCRIPT}; {IMAGE}, {QUESTION}, {OPTIONS}; latest affirmative message {AFF_ANS}; {OPP_ANS}. Instruction: Disagree when warranted. Provide an alternative strictly from {OPTIONS} and justify using visual evidence or logic. Output: A counter-argument plus one choice from {OPTIONS}.
Judge (per round) Receives: {TRANSCRIPT}; {IMAGE}, {QUESTION}, {OPTIONS}; {AFF_ANS}, {NEG_ANS}. Instruction: Evaluate (1) accuracy w.r.t. image and question and (2) justification quality. Decide if there is a clear preference and provide a tentative answer if any. Output (JSON only): {”Preference”:”Yes—No”,”Supported Side”:”Affirmative—Negative”,”Reason”:”…”,”debate_answer”: one of {OPTIONS} or ””}.
Judge (Finalization) Receives: {TRANSCRIPT}; {IMAGE}, {QUESTION}, {OPTIONS}; possibly reduced {OPTIONS2}. Instruction: (A) List candidate answers strictly from {OPTIONS}. (B) Provide the final decision with a brief reason. Output (JSON): {”Reason”:”…”,”debate_answer”: one of {OPTIONS2}}.
Table 5: List of extracted 41 features for the debate decision classifier.
Source Category Feature Name Description
Question Surface-level statistics QuestionToken Number of tokens in the question
AnswerToken Number of tokens in the model’s final answer
question_named_entity_count Number of named entities (e.g., person, location, date)
Uncertainty-related lexical cues qtype_what Indicator if the question begins with “what”
qtype_where Indicator if the question begins with “where”
qtype_why Indicator if the question begins with “why”
qtype_how Indicator if the question begins with “how”
qtype_when Indicator if the question begins with “when”
qtype_who Indicator if the question begins with “who”
qtype_is Indicator if the question begins with “is”
qtype_are Indicator if the question begins with “are”
qtype_does Indicator if the question begins with “does”
qtype_do Indicator if the question begins with “do”
qtype_other Indicator for other question types
Syntactic features QSyntacticDepth Maximum syntactic parse tree depth of the question
Readability metrics Question_flesch_reading_ease Flesch Reading Ease score of the question
Question_coleman_liau_index Coleman-Liau Index score of the question
Part-of-speech counts Question_num_nouns Number of nouns in the question
Question_num_verbs Number of verbs in the question
Question_num_adjs Number of adjectives in the question
Initial Reasoning Uncertainty-related lexical cues InitialConfidence Model’s confidence score for the initial answer
InitialReason_HedgeCount Number of hedge words (e.g., “maybe”, “likely”)
InitialReason_CertaintyCount Number of certainty words (e.g., “definitely”, “certainly”)
InitialReason_contrast Number of contrastive discourse markers (e.g., “but”, “however”)
Syntactic features InitialReason_SyntacticDepth Max syntactic parse depth of the justification
Readability metrics InitialReason_flesch_reading_ease Flesch Reading Ease score of the justification
InitialReason_coleman_liau_index Coleman-Liau Index of the justification
Part-of-speech counts InitialReason_num_nouns Number of nouns in the justification
InitialReason_num_verbs Number of verbs in the justification
InitialReason_num_adjs Number of adjectives in the justification
Self- Critique Uncertainty-related lexical cues NegativeConfidence Model’s confidence score for the alternative answer
NegativeReason_HedgeCount Number of hedge words in the critique
NegativeReason_CertaintyCount Number of certainty words in the critique
NegativeReason_contrast Number of contrastive discourse markers in the critique
FinalConfidence Model’s confidence score for the final answer
Syntactic features NegativeReason_SyntacticDepth Max syntactic parse depth of the critique
Readability metrics NegativeReason_flesch_reading_ease Flesch Reading Ease score of the critique
NegativeReason_coleman_liau_index Coleman-Liau Index of the critique
Part-of-speech counts NegativeReason_num_nouns Number of nouns in the critique
NegativeReason_num_verbs Number of verbs in the critique
NegativeReason_num_adjs Number of adjectives in the critique
A.2 Extracted Features for Debate Decision Classification
In this section, we provide a detailed description of the features that drive the debate decision classifier and evaluate their importance and necessity.
Feature Details.
[…]
To provide more details on the feature set, we summarize all extracted 41 features and their descriptions in Table 5, organized into five categories: (i) surface-level statistics (3 features from the input question only) such as token counts, which quantify response verbosity and completeness, as shorter or underdeveloped responses often correlate with lower reasoning quality; (ii) readability metrics (the input question, initial reasoning, and self-critique each with 2 features) like Flesch Reading Ease (davenport2014readabilitytweetsgeographiccorrelation) and Coleman-Liau Index (coleman), which assess how easy or complex a response is to read and potentially reveal vague or oversimplified reasoning; (iii) syntactic features (the input question, initial reasoning, and self-critique each with 1 features) such as parse tree depth, which indicate grammatical complexity and elaboration in reasoning structure; (iv) part-of-speech counts (the input question, initial reasoning, and self-critique each with 3 features) including noun, verb, and adjective frequencies, which help characterize the factual density, action-oriented phrasing, and descriptive richness of the response; and (v) uncertainty-related lexical cues (11 features from the input question, 4 features from the initial reasoning, and 5 features from the self-critique) that reflect reasoning quality, such as hedging phrases (e.g., “maybe”), certainty expressions (e.g., “definitely”), and contrastive connectives (e.g., “however” and “but”), which are commonly associated with hesitation or internal conflict in reasoning.
These features capture fine-grained characteristics of the response that relate to model confidence, uncertainty, and reasoning behavior, enabling the classifier to determine when MAD is likely to be beneficial based on the single-agent output. Since all features are derived directly from this structured output without requiring additional queries or repeated sampling, the classifier can make efficient and accurate decisions while preserving scalability and improving token efficiency.
Feature Importance with SHAP and PCA.
To understand which signals drive the debate decision, we evaluate two complementary metrics on the 41 features. SHAP importance (SHAP) quantifies each feature’s contribution to the classifier’s predicted score on each example. We report the mean absolute value of SHAP (typical range 0.01 to 0.5) in Fig. 3(a), where larger bars indicate stronger influence on decisions. PCA contributions (pca) capture global structure by aggregating feature loadings on principal components that explain most of the variance; larger bars indicate features that account for more variance (typical range 0.02 to 0.15 for each feature), as shown in Fig. 3(b). In both figures, signals from the self critique and the response structure rank near the top: uncertainty and conflict markers, length and brevity cues, and syntactic depth are consistently among the most influential features. The consistency between SHAP and PCA suggests that the classifier relies on stable, human understandable cues.
Feature Necessity.
To further investigate whether low-ranked features are still useful, we normalize the absolute SHAP values and PCA contributions for each feature, sum them to obtain a joint importance score for every feature, exclude the 20% of features with the lowest joint scores, and rerun the classifier. Concretely, we drop these eight features that consistently appear in the bottom 20%: qtype_do, qtype_who, qtype_when, qtype_how, qtype_where, qtype_is, qtype_does, and NegativeReason_CertaintyCount. Table 6 reports accuracy and tokens per dataset for iMAD with all features compared to the set that excludes these least important 20% features. As shown in Table 6, removing the low-ranked features leads to a small accuracy drop of about 0.5% and a noticeable increase in tokens of about 6.9% on average across six datasets, from 2,053 to 2,194 tokens per question. The higher token usage comes from less effective decisions that trigger debate on cases that are already correct, while missing a few recoverable hard cases mainly affects accuracy. These results justify that even features with lower importance scores in terms of SHAP and PCA rankings still contribute to more accurate debate decisions and help maintain token efficiency. Therefore, we use all 41 features.
Dataset iMAD (w/o Bottom 20%) iMAD (All Features)
Acc (%) # Token Acc (%) # Token
MEDQA 81.7 1,381 82.0 1,300
MMLU 88.7 1,061 89.2 1,010
GSM8K 84.0 1,107 84.8 1,025
OKVQA 90.0 2,783 90.3 2,601
VQA-v2 80.4 3,633 81.3 3,489
ScienceQA 90.8 3,196 90.8 2,893
Average 85.9 2,194 86.4 2,053
Table 6: Feature necessity study in iMAD. We normalize the SHAP importance scores and PCA loadings for each feature, sum them to obtain a joint importance score, and remove the bottom 20% of features. It shows an accuracy drop of about 0.5 % with the higher token usage of about 6.9% per question on average, showing the importance of all 41 features.
(a) SHAP Importance
(b) PCA Contributions
Figure 3: Feature importance for the debate-decision classifier. Bars indicate mean absolute SHAP values (top) and aggregated PCA contributions across principal components that explain most variance (bottom).
Appendix B Additional Analysis
In this section, we provide additional performance analysis. In section B.1, we analyze accuracy per 100k total tokens (ApT). In section B.2, we analyze per-question latency and explain how the mix of input and output tokens affects total inference time.
B.1 Analysis on Accuracy per Token
To further evaluate token efficiency, we report accuracy per tokens (ApT), where tokens include both inputs and outputs per instance. For each method and dataset, we first measure the average number of token tokens per question, summing both input and output tokens. Given accuracy $Acc\in[0,1]$ and the average total token count $\text{Token}$ (see Table 3 for exact values), we define
$\mathrm{ApT}=Acc\cdot\frac{100,000}{\#~\text{Token}}.$
A larger ApT indicates better token efficiency, corresponding to more correct answers per 100,000 tokens of budget.
Table 7 summarizes ApT across all datasets. CoT attains the highest ApT because it uses very few tokens per question, but the Table 3 shows that its prediction accuracy is consistently lower than that of multi-agent methods. SC improves accuracy over CoT by running CoT multiple times and aggregating the outputs, but the additional generations make its ApT much smaller, since the token cost increases faster than the accuracy gain. Full-debate MAD baselines such as MAD and GD achieve higher accuracy but pay a high token cost, which leads to very low ApT. Selective debate methods strike a better balance. Both DOWN and iMAD obtain much higher ApT than MAD and GD because they skip debating on many questions that the single agent already answers correctly. Compared with DOWN, iMAD has slightly lower ApT, about $0.92\times$ on average across the six datasets, because it still triggers debates on questions where debate is necessary to improve the answer. These debates consume many tokens, so ApT becomes slightly lower even though accuracy is higher. This modest extra token cost yields an average accuracy gain of about 4.1% without any threshold tuning on the evaluation datasets. In practice, this tradeoff is often favorable: iMAD retains most of the token savings of selective debate while achieving higher accuracy consistently across datasets. Thus, Table 7 should be read together with the accuracy table Table 3. ApT shows that iMAD preserves strong token efficiency compared with full-debate MAD and single-agent baselines, while the accuracy results highlight its consistent gains over DOWN.
Dataset Single-Agent
Full-Debate
MAD
Selective
MAD
CoT SC MAD GD DOWN iMAD
MEDQA 117.3 22.2 20.3 4.8 68.2 63.1
MMLU 113.6 23.4 26.7 6.2 98.0 88.3
GSM8K 115.4 20.6 22.2 4.8 89.4 82.7
OKVQA 45.4 8.1 11.5 2.6 37.6 34.7
VQA-v2 34.5 5.5 9.2 2.3 24.1 23.3
ScienceQA 50.0 8.8 13.2 3.2 34.5 31.4
Average 79.4 14.8 17.2 4.0 58.6 53.9
Table 7: Accuracy per 100k tokens (ApT). ApT measures the expected number of correct answers obtained for every 100,000 tokens, so larger values indicate better token efficiency.
Dataset Single-Agent Full-Debate Selective MAD
(s) MAD (s) (s)
CoT SC MAD GD DOWN iMAD
MEDQA 1.4 9.4 11.5 49.2 1.8 1.8
MMLU 1.0 5.3 5.1 29.7 1.1 1.1
GSM8K 1.6 8.7 8.0 29.8 1.6 1.6
OKVQA 1.4 12.9 9.1 37.1 1.7 1.7
VQA-v2 1.3 14.9 8.1 33.6 1.5 1.6
ScienceQA 1.5 15.2 10.9 44.3 1.8 1.8
Table 8: Inference time (s) per question comparison of iMAD and baselines across datasets.
Single-Agent Full-Debate MAD Selective MAD
Dataset CoT SC MAD GD DOWN iMAD
In Out In Out In Out In Out In Out In Out
MEDQA 478 175 2,307 1,175 2,834 1,200 10,682 6,150 981 180 1,110 190
MMLU 610 154 2,956 816 2,565 783 8,642 4,574 732 169 840 170
GSM8K 415 203 2,518 1,104 2,635 811 11,542 3,779 609 203 822 203
OKVQA 1,865 80 10,294 737 7,463 340 31,812 2,120 2,154 190 2,396 205
VQA-v2 2,175 70 13,211 802 8,346 450 34,282 1,809 3,077 185 3,291 198
ScienceQA 1,635 85 8,972 861 6,290 487 24,414 2,510 2,329 190 2,690 203
Table 9: The number of Input (In) tokens and output (Out) tokens per question comparison of iMAD and baselines across datasets. The sum of input token and output token equals the total # Token reported in Table 3.
B.2 Inference Time Analysis
To evaluate the computational efficiency of iMAD, we measure the per-question inference time across different methods. As shown in Table 8, CoT achieves the lowest inference latency but yields lower accuracy. SC improves accuracy by running CoT multiple times and aggregating the outputs, but this significantly increases latency (up to 5.3–15.2s). Full-debate MAD frameworks such as MAD and GD incur even higher latency, with inference times reaching up to 49.2s per question. In contrast, selective MAD methods such as DOWN and iMAD maintain near single-agent latency, with inference times ranging from 1.1 to 1.8s. However, unlike DOWN, which suffers from lower accuracy due to its inability to generalize across different models or datasets, iMAD consistently matches or exceeds the accuracy of full-debate MAD frameworks. On average, iMAD reduces inference time by 32.6% and 68.9% compared to MAD and GD, respectively, while preserving low latency. It is worth noting that latency is not directly proportional to the total token cost. Table 3 reports total tokens (input and output), and Table 9 decomposes them into input and output (out). Output tokens generally take longer per token to generate than input tokens, although very large inputs still add measurable time. For example, GSM8K uses fewer total tokens than MMLU (618 vs 764), yet produces more output tokens on average (203 vs 154), so decoding takes longer (1.6s vs 1.0s). Within a dataset, SC on VQA-v2 uses fewer output tokens than MAD (802 vs 2,450) but a much larger input (13,211 vs 6,346) because it aggregates multiple single-agent runs, which explains its higher latency despite lower outputs. Overall, these findings highlight the ability of iMAD to balance accuracy and computational cost effectively, making it a practical choice for real-world deployments where both accuracy and inference speed are critical.
Appendix C Additional Results
In this section, we present two ablation studies and a cross-LLM evaluation to assess the effectiveness of iMAD ’s two key design components and to gauge its generality. The ablations examine (i) the structured self-critique prompting, designed to enable interpretable and informative feature extraction for downstream debate decision classification (in section C.1); and (ii) the FocusCal loss, designed to guide the classifier in skipping debates that are unnecessary (i.e., when the single-agent answer is already correct or not recoverable by MAD) and triggering them when beneficial (i.e., when MAD is likely to correct an initially wrong answer) (in section C.2). In section C.3, We present cross-LLM results on GPT-5 nano and Qwen 3.0 models.
C.1 Ablation Studies: Effectiveness of Structured Self-Critique Prompting
To evaluate the impact of prompt design, we compare structured self-critique prompt in iMAD, which asks the model to first provide reason step by step and then explicitly critique its own answer, against a standard CoT prompt without self-critique. For this study, we report (i) input-output token usage per data instance and (ii) accuracy of the answer (i.e., final answer correctness), to assess the overall impact of the prompt on the performance of system. As shown in Table 10, ours incorporating self-critique consistently improves accuracy across all six datasets. This improvement is particularly notable on complex reasoning tasks; for example, GSM8K shows a 7.2% accuracy gain. Although the self-critique prompt slightly increases token usage (e.g., from 1,222 to 1,300 on MEDQA), this cost is justified by the substantial accuracy improvement.
This accuracy improvement comes from the structured self-critique prompt’s ability to elicit more detailed and diverse reasoning, including counterarguments and comparative confidence judgments. These elements help reveal subtle indicators of uncertainty or conflicting logic in the model’s response, which are often missing from standard CoT outputs. As a result, the debate decision classifier can more effectively determine whether a debate is warranted. For instance, the structured prompt may generate divergent reasoning paths or express comparable confidence across competing answers. The classifier learns to interpret such cues as signs of uncertainty, indicating that debate may improve the final answer. In contrast, standard CoT prompts typically produce less informative responses, which makes it more difficult to identify cases where debate could be beneficial. By leveraging lexical and semantic features such as contrastive phrasing, shifts in certainty, and justification depth, the 41 extracted features help the classifier make more accurate debate decisions.
Dataset Prompt w/o Self-Critique Prompt w/ Self-Critique
Acc (%) # Token Acc (%) # Token
MEDQA 80.2 1,222 82.0 1,300
MMLU 88.3 943 89.2 1,010
GSM8K 77.6 988 84.8 1,025
OKVQA 88.1 2,342 90.3 2,601
VQA-v2 78.9 3,314 81.3 3,489
ScienceQA 88.4 2,541 90.8 2,893
Table 10: Performance of single-agent prompts with and without self-critique across QA and VQA datasets.
Dataset BCE (%) MSE (%)
FocusCal (%)
MEDQA 84.2 83.1 88.8
MMLU 85.0 86.7 92.4
GSM8K 88.2 87.9 91.4
OKVQA 89.3 89.1 95.9
VQA-v2 89.2 89.9 92.8
ScienceQA 91.4 91.8 93.4
Table 11: Accuracy of beneficial debate decisions using BCE or MSE compared to our FocusCal Loss.
C.2 Ablation Study: Effectiveness of FocusCal Loss
To evaluate the effectiveness of our proposed FocusCal loss, we conduct the evaluation in three aspects: (i) how each loss component of $L_{\text{AF}}$, $L_{\text{CP}}$, and ECE in our FocusCal contributes to the final performance in terms of accuracy and token costs; (ii) how FocusCal outperforms standard losses such as Binary Cross-Entropy (BCE) and Mean Squared Error (MSE) on debate decision correctness; and (iii) what is the impact of the calibration term ECE in FocusCal while replacing only ECE with BCE (li2024rediscoveringbcelossuniform) or MSE (koksoy2006multiresponse) while keeping $L_{\text{AF}}$ and $L_{\text{CP}}$ fixed.
Loss Component Acc (%) # Token
$L_{\text{AF}}$ only 78.8 3,558
$L_{\text{CP}}$ only 78.1 3,379
ECE only 79.1 3,757
$L_{\text{AF}}$ + $L_{\text{CP}}$ 79.7 3,552
$L_{\text{AF}}$ + ECE 79.7 3,561
$L_{\text{CP}}$ + ECE 79.8 3,577
FocusCal ($L_{\text{AF}}$ + $L_{\text{CP}}$ + ECE) 81.3 3,489
Table 12: Ablation study on three loss terms in FocusCal using the VQA-v2 dataset in terms of accuracy (Acc) and total token cost (# Token).
FocusCal Loss Terms.
In Table 12, we report accuracy (Acc) and total token cost (# Token) on the VQA-v2 dataset using subsets of the three components in FocusCal (i.e., $L_{\text{AF}}$, $L_{\text{CP}}$, and ECE). We mainly use the VQA-v2 dataset for brevity since the other datasets follow the same performance pattern. When using each single loss term, the behaviors are complementary: $L_{\text{CP}}$ yields the fewest tokens, 3,379 tokens in total, but the lowest accuracy of 78.1% by penalizing overconfident scores raises more examples above the decision threshold, causing the classifier to skip debates, including recoverable ones. ECE attains the best single-term accuracy of up to 79.1% but with the highest token cost of 3,757, as calibration makes scores more faithful to uncertainty and flags more borderline examples for debate, increasing debate frequency even when some debates are unnecessary. $L_{\text{AF}}$ improves accuracy over $L_{\text{CP}}$ by emphasizing hard negatives and lowering scores for wrong-but-confident answers, which triggers helpful debates on those instances.
Two-term combinations consistently improve over single terms. On VQA-v2, using $L_{\text{AF}}~and~\text{ECE}$ and using $L_{\text{CP}}~and~\text{ECE}$ raise accuracy to 79.7% and 79.8% with 3,561 and 3,577 tokens, respectively. This is because ECE term encourages the predicted debate-triggering score to align with empirical correctness, while $L_{\text{AF}}$ penalizes overconfident errors and emphasizes borderline cases, and $L_{\text{CP}}$ penalizes misalignment between the model’s confidence and the semantic uncertainty expressed in the response. In comparison, using $L_{\text{AF}}~and~L_{\text{CP}}$ attains 79.7% with 3,552 tokens. Without ECE, the debate-triggering score is less aligned with empirical correctness, so accuracy is similar but tokens are slightly lower. Using all three terms in FocusCal achieves the highest accuracy (up to 81.3%) with fewer tokens (3,489 tokens), since ECE aligns the debate-triggering score with empirical correctness, $L_{\text{AF}}$ penalizes overconfident errors and emphasizes borderline cases, and $L_{\text{CP}}$ penalizes misalignment between confidence and semantic uncertainty. All three terms together help to better identify recoverable errors and avoid unnecessary debates.
FocusCal vs. BCE/MSE.
We compare our FocusCal loss with two commonly used loss functions to train the binary debate decision classifiers: (i) BCE and (ii) MSE. We measure the percentage of beneficial decisions made by the classifier, defined as correctly skipping debate for correct answers or unrecoverable errors, and correctly triggering debate for recoverable errors. As shown in Table 11, FocusCal achieves the highest rate of beneficial debate decisions across all six datasets, reaching 95.9% on OKVQA and substantially outperforming both BCE (89.3%) and MSE (89.1%). The improvement in beneficial debate decision accuracy stems from FocusCal’s targeted design, which enables the classifier to address two key issues: overconfident incorrect predictions and misalignment between predicted scores and uncertainty signals revealed in the single-agent response. These issues are not effectively handled by BCE or MSE, because they penalize all classification errors equally.
Dataset BCE MSE ECE
Acc (%) # Token Acc (%) # Token Acc (%) # Token
MEDQA 81.2 1,508 81.4 1,389 82.0 1,300
MMLU 89.0 1,296 88.8 1,005 89.2 1,010
GSM8K 83.9 1,115 82.5 1,099 84.8 1,025
OKVQA 89.5 2,887 89.1 2,588 90.3 2,601
VQA-v2 80.7 3,626 79.8 3,798 81.3 3,489
ScienceQA 90.8 3,352 89.4 2,899 90.8 2,893
Table 13: Accuracy (Acc) and average token cost (# Token) per question when replacing the ECE term in FocusCal with BCE or MSE while maintaining $L_{\text{AF}}$ and $L_{\text{CP}}$.
Calibration Term of ECE Compared to MSE and BCE.
To further evaluate the design choice of the calibration loss component, we fix $L_{\text{AF}}$ and $L_{\text{CP}}$ and replace the ECE term in FocusCal with BCE or MSE. As shown in Table 13, ECE attains the best or tied accuracy on all datasets and usually uses fewer tokens on average. For example, on GSM8K, ECE improves accuracy over MSE from 82.5% to 84.8% while using about 7% fewer tokens, reducing token usage from 1,099 to 1,025. On VQA-v2, it increases accuracy over BCE from 80.7% to 81.3% and lowers token usage by about 4%, from 3,626 to 3,489. On MMLU, ECE uses a similar number of tokens as MSE (1,010 vs 1,005), but attains higher accuracy, and compared with BCE it improves accuracy while cutting token usage by roughly 22%, from 1,296 to 1,010. Averaged across six datasets, ECE yields 86.4% accuracy with 2,053 tokens, versus 85.85% and 2,297 for BCE, and 85.17% and 2,130 for MSE. These results indicate that ECE calibrates better, whereas BCE and MSE treat errors uniformly and provide weaker calibration, causing borderline cases to either trigger unnecessary debates or skip debates that could recover errors.
C.3 Additional Cross-LLM Evaluation: Qwen 3.0 and GPT-5 nano
Single-Agent Full-Debate MAD Selective MAD
Dataset CoT SC MAD GD DOWN iMAD
Acc (%) # Token Acc (%) # Token Acc (%) # Token Acc (%) # Token Acc (%) # Token Acc (%) # Token
MEDQA 87.7 1,603 88.4 8,417 91.3 46,352 90.4 160,232 89.0 2,084 91.6 2,165
MMLU 85.1 1,482 86.5 8,351 87.7 40,224 87.8 142,784 86.0 1,927 87.8 2,071
GSM8K 70.4 1,550 73.6 8,925 75.9 38,500 74.9 134,750 74.8 2,015 76.5 2,280
OKVQA 88.0 1,750 89.0 9,225 89.2 62,000 89.3 215,000 88.5 2,275 89.8 2,400
VQA-v2 76.7 1,661 77.2 9,536 78.3 41,119 79.0 144,917 78.0 2,159 80.1 2,258
ScienceQA 86.8 1,600 86.8 8,002 88.5 59,000 88.0 201,500 87.2 2,080 89.4 2,260
Table 14: Performance on GPT-5 nano: Accuracy (Acc) and average token cost (# Token) per question.
Single-Agent Full-Debate MAD Selective MAD
Dataset CoT SC MAD GD DOWN iMAD
Acc (%) # Token Acc (%) # Token Acc (%) # Token Acc (%) # Token Acc (%) # Token Acc (%) # Token
MEDQA 80.3 1,030 81.0 5,465 83.2 178,063 82.5 311,739 81.5 1,339 83.8 1,551
MMLU 87.8 1,131 89.2 6,921 91.3 149,686 91.3 309,309 88.9 1,470 91.7 1,523
GSM8K 74.2 1,099 75.0 5,290 76.8 120,000 76.8 284,000 75.5 1,274 77.6 1,366
OKVQA 88.5 1,000 89.4 5,400 88.5 130,000 89.3 246,000 88.9 1,300 90.5 1,501
VQA-v2 76.5 912 76.9 5,116 78.2 81,618 79.0 159,560 76.8 1,186 79.4 1,350
ScienceQA 87.0 1.022 87.2 5,345 90.5 115,000 89.0 283,000 87.5 1,287 90.9 1,338
Table 15: Performance on Qwen 3.0: Accuracy (Acc) and average token cost (# Token) per question.
We evaluate iMAD on two additional base models, GPT-5 nano (gpt5) and Qwen 3.0 (yang2025qwen3technicalreport), with results in Tables 14 and 15, respectively. All prompt designs, the self-critique stage, memory synchronization, and the debate decision policy are the same as those in the main experimental settings (see Section 5) on Gemini 2.0 Flash. We conducted a one-time sanity check of prompts and thresholds on auxiliary datasets (PubMedQA and GQA) and then froze the debate decision classifier and its threshold. We report both accuracy and the average token cost per question for performance evaluation.
Across both base models, iMAD attains the highest accuracy on every dataset while holding token budgets close to DOWN and far below full-debate methods. On GPT-5 nano with VQA-v2, iMAD reaches 80.1% at 2,258 tokens; GD is 79.0% at 144,917 tokens and MAD is 78.3% at 41,119 tokens. Thus, iMAD matches or exceeds their accuracy while using 98.4% and 94.5% fewer tokens, respectively. On Qwen 3.0 with MMLU, iMAD reaches 91.7% at 1,523 tokens, whereas MAD and GD both achieve 91.3% accuracy but require 149,686 and 309,309 tokens. These correspond to 98.98% and 99.51% lower token cost for iMAD. Full-debate methods incur extremely high token costs because, in every round, they rebroadcast the growing transcript to all agents, expanding the input context even when the example is not recoverable. By contrast, iMAD triggers debate only when the classifier detects hesitation or conflict in the self-critique single-agent output, so extra tokens are spent only on incorrect cases that are likely to be corrected through debate. Moreover, as Tables 14 and 15 show, on reasoning-heavy tasks such as GSM8K, iMAD outperforms single-agent baselines on both models while adding only tens to a few hundred tokens per question. This is because self-critique cues identify cases where triggering debate can fix some errors in intermediate reasoning steps. By contrast, for vision or short factual tasks such as OKVQA, VQA-v2, and MMLU, accuracy gains are smaller, since when the decisive visual cue or fact is missing, triggering additional debate cannot introduce new evidence to flip the incorrect answer.
Compared to DOWN, iMAD achieves higher accuracy with only a small increase in token costs. On GPT-5 nano, iMAD raises MEDQA from 89.0% to 91.6% with 81 additional tokens per question (about 3.9%), GSM8K from 74.8% to 76.5% with 265 more tokens (about 13.2%), and ScienceQA from 87.2% to 89.4% with 180 more tokens (about 8.7%). On Qwen 3.0, iMAD raises MEDQA from 81.5% to 83.8% with 212 more tokens (about 15.8%), GSM8K from 75.5% to 77.6% with 92 more tokens (about 7.2%), and ScienceQA from 87.5% to 90.9% with 51 more tokens (about 4.0%). These gains arise because DOWN relies on the model’s raw confidence, which can be miscalibrated and overconfident, so it often skips debates that are actually needed. iMAD instead utilizes the self-critique responses and uses a FocusCal-trained debate classifier to detect hesitation and conflict, only spend extra tokens on the cases where triggering debate can correct the answer. These accuracy gains arise because DOWN relies on raw model confidence score, which is often miscalibrated and overconfident, leading it to skip debates that are actually needed. In contrast, iMAD first extracts structured features from the self-critique response and then uses a trained debate classifier to decide whether to trigger debate.
With identical prompts and decoding settings, Qwen 3.0 typically generates briefer single-agent responses than GPT-5 nano. This results in lower token budgets in iMAD while preserving similar accuracy gains. Overall, the debate-decision classifier trained once transfers across base models, delivering consistent accuracy improvements with a small token cost while avoiding the heavy cost of full debate.
Appendix D Discussion
While iMAD demonstrates strong improvements in both accuracy and token efficiency, here we discuss its design assumptions and potential extension.
Limited Labeled Data.
One key requirement for iMAD is constructing a training dataset with correctness labels for single-agent answers, which we create by prompting a single agent on two representative datasets (i.e., one QA and one VQA) and manually labeling each response as correct or incorrect. These binary correctness labels serve as supervision for training the debate decision classifier: each example is paired with features extracted from the self-critique response, and the classifier learns to predict whether a debate would be beneficial. Although this labeling process is lightweight and limited to only two datasets, our careful dataset selection exposes the classifier to diverse reasoning behaviors. This design allows the classifier to generalize well and perform effectively on six held-out datasets without further tuning. However, because the classifier is trained offline and remains fixed during deployment, it may not effectively adapt to changes in model behavior. Future work could explore online adaptation through feedback-driven updates or weak supervision to further improve generalization.
Prompt Engineering.
Another consideration lies in the reliance on structured self-critique prompting, which iMAD uses to extract interpretable features from the model’s output to train the downstream debate classifier to capture general model behavior. However, this design assumes that the underlying LLM agent can effectively articulate its reasoning and express uncertainty. In domains where the model’s responses are incoherent or fail to convey uncertainty clearly, the extracted features may become less informative. In such cases, additional prompt refinement may be necessary to preserve debate decision quality.
Future Work.
An interesting extension is to move the debate–triggering decision inside the generation process rather than waiting for the full self-critique to complete. In our current deployment, the LLM agent is accessed only through a black-box API (e.g., Gemini or GPT), so iMAD can decide whether to trigger debate only after the complete self-critique is returned. With open-source models or streaming APIs, we can instrument token-by-token generation, monitor the emerging rationale for early hesitation cues, and trigger MAD before the self-critique finishes, potentially reducing both latency and token costs while maintaining accuracy. A further direction is to combine these in-generation signals with internal model states (such as logit trajectories or entropy over the candidate answers) and learn an adaptive policy (e.g., reinforcement learning), that allocates debate budget dynamically over the course of generation.